Given a circle with a radius of 16 mm, find:
The area of the circle

The arch length for a central angle of θ=300°

Find the area of the sector created by a central angle of θ=300°

Answer by TimothyLamb(4379)

area = pi * r^2 = 3.1415 * (16*16) = 804.24772 sq.mm
---
arc length = 300/360 * 2 * pi * r = 83.775804 mm
---
area of sector = 300/360 * area = 670.20643 sq.mm
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php